# Dévissage for generation in derived categories

### v4

@@ 2. Generation
    $E$ is supported on $Z$ whenever $\operatorname{Supp}(E)$ is contained in $Z$
@@
    $D^{b}_{\operatorname{coh},Z}(X)$ is the strictly full subcategory of $D^{b}_{\operatorname{coh}}(X)$ consisting of objects whose support is contained in $Z$.
3. Results
This section will prove our main results and to start, we need something that is akin to generic freeness of coherent $\mathcal{O}_{X}$-modules on a Noetherian integral scheme. The following might be known to a few, but we offer an alternative proof for the sake of interest in our setting.
@@ 3. Results
Let $X$ be a Noetherian integral scheme. If $E$ and $G$ are objects of $D^{b}_{\operatorname{coh}}(X)$ with full support, then there exists an open immersion $i\colon U\to X$ such that $i^{\ast}E$ is in $\langle i^{\ast}G\rangle_{1}$ and $i^{\ast}G$ is in $\langle i^{\ast}E\rangle_{1}$.
@@
Consider a nonempty open affine $U$ in $X$. The set $S$ of points $p$ in $U$ such that $E_{p}$ is in $\langle G_{p}\rangle_{1}$ is Zariski open, see [Let21, Proposition 3.5]. Clearly, $S$ is nonempty as it contains the generic point of $X$. By the same reasoning, the set $T$ of points $p$ in $X$ such that $G_{p}$ is in $\langle E_{p}\rangle_{1}$ is Zariski open and nonempty. Choose a nonempty open affine $W$ in $S\cap T$. Let $j\colon W\to X$ be the associated open immersion. It follows that $j^{\ast}E$ is in $\langle j^{\ast}G\rangle_{1}$ and $j^{\ast}G$ is in $\langle j^{\ast}E\rangle_{1}$ because for each $p$ in $W$, one has $E_{p}$ is in $\langle G_{p}\rangle_{1}$ and $G_{p}$ is in $\langle E_{p}\rangle_{1}$, see [Let21, Corollary 3.4]. ∎
Remark 3.2.
A variation of Lemma 3.1 has appeared in the proof of [Rou08, Proposition 7.16] for separated reduced schemes of finite type over a field.
Theorem 3.3.
Let $X$ be a Noetherian scheme. If $P$ is a perfect complex on $X$ with full support, then $D_{\operatorname{coh}}^{b}(X)$ is classically generated by $P$ and objects of the form $i_{\ast}i^{\ast}P$ where $i\colon Z\to X$ is a closed immersion from an integral scheme.
Proof.
Let $\mathcal{S}$ be the collection consisting of $P$ and objects of the form $i_{\ast}i^{\ast}P$ where $i\colon Z\to X$ is a closed immersion from an integral scheme. If $X=\emptyset$, then there is nothing to check, so without loss of generality, $X$ is nonempty. We argue by Noetherian induction. Without loss of generality, assume the claim holds for any properly contained closed subscheme $Z$ of $X$; it is possible to reduce to this case by [Rou08, Lemma 7.40]. We may further impose that $X$ is reduced, see [Rou08, Lemma 7.40] or [Lan24, Remark 2.23]. Let $E$ be an object of $D^{b}_{\operatorname{coh}}(X)$.
If $E$ has support properly contained in $X$, then the induction hypothesis tells us that $E$ belongs to $\mathcal{S}$. Indeed, let $Z$ be the reduced induced closed subscheme structure on $\operatorname{supp}(Z)$ with corresponding ideal sheaf $\mathcal{I}$. There is an integer $n$ such that $\mathcal{I}^{n}E=(0)$, which ensures $E$ is isomorphic to an object of the form $i_{\ast}E^{\prime}$ for $E^{\prime}$ in $D^{b}_{\operatorname{coh}}(Z_{n})$ and $i\colon Z_{n}\to X$ the closed immersion associated to the ideal sheaf $\mathcal{I}^{n}$, see [Rou08, Lemma 7.40] or [Lan24, Remark 2.23]. Note that $i^{\ast}P$ is a classical generator for $\operatorname{perf}(Z_{n})$, see [Sta23, Tag 0BQT]. The induction hypothesis ensures that $i_{\ast}D^{b}_{\operatorname{coh}}(Z_{n})$ belongs to $i_{\ast}\mathcal{S}^{\prime}$ because $D^{b}_{\operatorname{coh}}(Z_{n})=\mathcal{S}^{\prime}$ where $\mathcal{S}^{\prime}$ is the collection consisting of $i^{\ast}P$ and objects of the form $t_{\ast}(i\circ t)^{\ast}P$ where $t\colon Z^{\prime}\to Z_{n}$ is a closed immersion from an integral scheme. Note that $i_{\ast}\mathcal{S}^{\prime}$ is contained in $\mathcal{S}$, and so, it follows that $E$ belongs to $\langle\mathcal{S}\rangle$ as claimed.
We will now consider the case where $E$ has full support. Moreover, as objects in $D^{b}_{\operatorname{coh}}(X)$ are finitely built by their cohomology sheaves, we can impose that $E$ is a coherent $\mathcal{O}_{X}$-module. First, lets consider the case where $X$ is integral. We have that $E$ is a coherent $\mathcal{O}_{X}$-module with full support on the integral scheme $X$. There exists a dense open immersion $j\colon U\to X$ such that $j^{\ast}E$ and $j^{\ast}P$ finitely build one another in at most one cone in $D^{b}_{\operatorname{coh}}(U)$, see Lemma 3.1. If $U=X$, then we are done, so assume $U$ is a properly contained subset of $X$. Note that $j^{\ast}E$ is a direct summand of an object of the form $\oplus_{n\in\mathbb{Z}}j^{\ast}P^{\oplus r_{n}}[n]$, so we have an object $j^{\ast}E$ and isomorphism $\phi\colon j^{\ast}A\oplus j^{\ast}E\to\oplus_{n\in\mathbb{Z}}j^{\ast}P^{\oplus r_{n}}[n]$ in $D^{b}_{\operatorname{coh}}(U)$. There exists a roof in $D^{b}_{\operatorname{coh}}(X)$:
${B}$${{\bigoplus_{n\in\mathbb{Z}}P^{\oplus r_{n}}[n]}}$${{A\oplus E}.}$$\scriptstyle{f}$$\scriptstyle{g}$
Observe that the cones of $f,g$ are supported on the properly contained closed subset $Z$ of $X$ as the map $j^{\ast}\colon D^{b}_{\operatorname{coh}}(X)\to D^{b}_{\operatorname{coh}}(U)$ is a Verdier localization. This ensures the objects $\operatorname{cone}(f),\operatorname{cone}(g)$ are in $\mathcal{S}$ by our inductive hypothesis. Therefore, $B$ belongs to $\mathcal{S}$, and hence, so does $A\oplus E$.
Lastly, we will consider the case where $X$ is not integral. Denote the irreducible components of $X=\operatorname{Supp}(E)$ by $Z_{1},\ldots,Z_{n}$. There exists a short exact sequence of coherent $\mathcal{O}_{X}$-modules:
$0\to E_{1}\to E\to E_{1}^{\prime}\to 0$
where $\operatorname{Supp}(E_{1})\subseteq Z_{1}$ and $\operatorname{Supp}(E_{1}^{\prime})\subseteq\cup^{n}_{i=2}Z_{i}$, see [Sta23, Tag 01YD]. Both objects $E_{1}$ and $E_{1}^{\prime}$ are supported on properly contained closed subsets of $X$, and so, each belong to $\langle\mathcal{S}\rangle$ according to the induction hypothesis. However, this implies that $E$ belongs to $\langle\mathcal{S}\rangle$, which completes the proof. ∎
Example 3.4.
Let $X$ be a Noetherian scheme. By Theorem 3.3, we have that $D^{b}_{\operatorname{coh}}(X)$ is classically generated by $\mathcal{O}_{X}$ and objects of the form $i_{\ast}\mathcal{O}_{Z}$ where $i\colon Z\to X$ is a closed immersion from an integral scheme.
Theorem 3.5.
Let $X$ be a Noetherian scheme. If $P$ is a classical generator for $\operatorname{perf}(X)$, then $D^{b}_{\operatorname{coh}}(X)$ is classically generated by $P$ and the objects of the form $i_{\ast}i^{\ast}P$ where $i\colon Z\to X$ is a closed immersion from an integral scheme such that $i(Z)$ is contained in $\operatorname{sing}(X)$.
Proof.
Denote by $\mathcal{U}$ for the collection of objects consisting of $P$ and those of the form $i_{\ast}i^{\ast}P$ where $i\colon Z\to X$ a closed immersion from an integral scheme such that $i(Z)$ is contained in $\operatorname{sing}(X)$. We will show that $\langle\mathcal{U}\rangle=D^{b}_{\operatorname{coh}}(X)$. Note that $\mathcal{O}_{W}$ is finitely built by $t^{\ast}P$ for each closed immersion $t\colon W\to X$ [Sta23, Tag 0BQT], and so, $t_{\ast}\mathcal{O}_{W}$ is finitely built by $t_{\ast}t^{\ast}P$. This ensures that $i_{\ast}\mathcal{O}_{Z}$ is contained in $\langle\mathcal{U}\rangle$ for each closed immersion from an integral scheme such that $i(Z)$ is contained in $\operatorname{sing}(X)$.
Consider an affine open cover $X=\bigcup^{n}_{i=1}U_{i}$ with $s_{i}\colon U_{i}\to X$ being the associated open immersions. Note that for each $1\leq i\leq n$, one has $\langle s^{\ast}_{i}\mathcal{U}\rangle$ containing all objects of the form $p_{\ast}\mathcal{O}_{W}$ where $p\colon W\to U_{i}$ is a closed immersion where $W$ such that $p(W)$ is contained in $\operatorname{sing}(U_{i})$, including the structure sheaf. It follows that $\langle s^{\ast}_{i}\mathcal{U}\rangle=D^{b}_{\operatorname{coh}}(U_{i})$ because $U_{i}$ is affine for each $1\leq i\leq n$, see [Tak14, Corollary 4.3.(1)].
Let $E$ be an object of $D^{b}_{\operatorname{coh}}(X)$. There is an object $G_{i}$ in $\langle\mathcal{U}\rangle_{1}$ such that $s^{\ast}_{i}E$ is finitely built by $s^{\ast}_{i}G_{i}$ in $D^{b}_{\operatorname{coh}}(U_{i})$ because $\langle s^{\ast}_{i}\mathcal{U}\rangle=D^{b}_{\operatorname{coh}}(U_{i})$. Set $G=\oplus^{n}_{j=1}G_{i}$. It follows that $s^{\ast}_{j}E$ is finitely built by $s^{\ast}_{j}G$ in $D^{b}_{\operatorname{coh}}(U_{j})$ for each $1\leq j\leq n$. This ensures that $E_{p}$ is finitely built by $G_{p}$ in $D^{b}_{\operatorname{coh}}(\mathcal{O}_{X,p})$ for each $p$ in $X$. Hence, $E_{p}\overset{\mathbb{L}}{\otimes}K(p)$ is finitely built by $G_{p}$ in $D^{b}_{\operatorname{coh}}(\mathcal{O}_{X,p})$ for each $p$ in $X$ where $K(p)$ is the Koszul complex on a minimal set of generators for the maximal ideal of the local ring $\mathcal{O}_{X,p}$. Indeed, $K(p)$ is finitely built by $\mathcal{O}_{X,p}$, and tensoring with $E_{p}$ tells us that $E_{p}\overset{\mathbb{L}}{\otimes}K(p)$ is finitely built by $E_{p}$.
If $E_{p}\overset{\mathbb{L}}{\otimes}K(p)$ is finitely built by $G_{p}$ in $D^{b}_{\operatorname{coh}}(\mathcal{O}_{X,p})$ for each $p$ in $X$, then $E$ is finitely built by $P\overset{\mathbb{L}}{\otimes}G$ in $D^{b}_{\operatorname{coh}}(X)$, see [BIL+23, Theorem 1.7]. As $G$ belongs to $\langle\mathcal{U}\rangle_{1}$, it is a direct summand of an object of the form $\oplus_{l\in\mathbb{Z}}(i_{l,\ast}i^{\ast}_{l}P)^{\oplus r_{l}}[l]$. By tensoring with $P$ and using projection formula, it follows that $P\overset{\mathbb{L}}{\otimes}\big{(}\oplus_{l\in\mathbb{Z}}(i_{l,\ast}i^{\ast}_{l}P)^{\oplus r_{l}}[l])$ is isomorphic to $\oplus_{l\in\mathbb{Z}}(i_{l,\ast}i^{\ast}_{l}(P\overset{\mathbb{L}}{\otimes}P))^{\oplus r_{l}}[l]$. Note that $P\overset{\mathbb{L}}{\otimes}G$ is a direct summand of $\oplus_{l\in\mathbb{Z}}(i_{l,\ast}i^{\ast}_{l}(P\overset{\mathbb{L}}{\otimes}P))^{\oplus r_{l}}[l]$. As $P\overset{\mathbb{L}}{\otimes}P$ is a classical generator for $\operatorname{perf}(X)$, we see that $\oplus_{l\in\mathbb{Z}}(i_{l,\ast}i^{\ast}_{l}(P\overset{\mathbb{L}}{\otimes}P))^{\oplus r_{l}}[l]$ is finitely built by $\oplus_{l\in\mathbb{Z}}(i_{l,\ast}i^{\ast}_{l}P)^{\oplus r_{l}}[l]$. Therefore, $E$ must be contained in $\langle\mathcal{U}\rangle$ because it is finitely built by $\oplus_{l\in\mathbb{Z}}(i_{l,\ast}i^{\ast}_{l}P)^{\oplus r_{l}}[l]$. ∎
Remark 3.6.
  1. (1)
    It is important to assume $P$ is a classical generator for $\operatorname{perf}(X)$ in Theorem 3.5. A possible relaxation might be imposing $P$ be a perfect complex on $X$ with full support, but we see that there are counterexamples, i.e. $P=\mathcal{O}_{X}$ where $X=\mathbb{P}^{1}_{k}$ for $k$ a field because $\mathcal{O}_{X}$ is not a strong generator for $D^{b}_{\operatorname{coh}}(\mathbb{P}^{1}_{k})$ and the singular locus is empty.
  2. (2)
    Theorem 3.5 has appeared in the different forms in the literature, but not in our generality. This includes the affine setting [Tak14, Corollary 4.3], proof for [Rou08, Theorem 7.38], and divisorial varieties [BFK14, Lemma 4.13].
A possible utility for Theorem 3.5 in practice is when the singular locus is finite, which we demonstrate a special case below.
Example 3.7.
Let $X$ be a normal irreducible quasi-projective surface over a field. The singular locus of $X$ must be finite, say with closed points $p_{1},\cdots,p_{n}$ and associated closed immersions $\pi_{i}\colon\operatorname{Spec}(\kappa(p_{i}))\to X$. If $\mathcal{L}$ is an ample line bundle on $X$, then $\mathcal{L}\oplus(\oplus^{n}_{i=1}\pi_{i,\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p_{i}))})$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$.
It is possible to show that $\langle\mathbb{R}\pi_{\ast}D^{b}_{\operatorname{coh}}(Y)\rangle=D^{b}_{\operatorname{coh}}(X)$ for $\pi\colon Y\to X$ a proper surjective morphism by Noetherian induction coupled with Theorem 3.5. We refine this to a stronger statement in what follows.
@@
Let $\pi\colon Y\to X$ be a proper surjective morphism of Noetherian schemes. For every object $K$ in $\mathbb{R}\pi_{\ast}D^{-}_{\operatorname{coh}}(Y)$ and integer $a$, there is a map $f\colon K\to L$ with $L$ in $\mathbb{R}\pi_{\ast}D^{b}_{\operatorname{coh}}(Y)$ such that $\operatorname{cone}(f)$ is in $D^{\leq a}_{\operatorname{Qcoh}}(X)$.
Proof.
Let $K:=\mathbb{R}\pi_{\ast}E$. Then $L=\mathbb{R}\pi_{\ast}(\tau_{\geq a-N}E)$ fits the bill for some integer $N$. ∎
Lemma 3.9.
Let $\pi\colon Y\to X$ be a proper surjective morphism of Noetherian schemes. Let $n\geq 0$. For each $K$ in $D^{-}_{\operatorname{coh}}(X)$ which belongs to $\big{(}\mathbb{R}\pi_{\ast}D^{-}_{\operatorname{coh}}(Y)\big{)}^{\star n}$ and integer $a$, there is a map $f\colon K\to L$ with $L$ an object of $\big{(}\mathbb{R}\pi_{\ast}D^{b}_{\operatorname{coh}}(Y)\big{)}^{\star n}$ such that $\operatorname{cone}(f)$ in $D^{<a}_{\operatorname{Qcoh}}(X)$.
@@
where $A,B$ are objects of $\mathbb{R}\pi_{\ast}D^{-}_{\operatorname{coh}}(Y)$. Then Lemma 3.8 ensures there is a commutative square:
[…]
@@
Let $\pi\colon Y\to X$ be a proper surjective morphism of Noetherian schemes. If $E$ is an object of $D^{b}_{\operatorname{coh}}(X)\cap\langle\mathbb{R}\pi_{\ast}D^{-}_{\operatorname{coh}}(Y)\rangle_{n}$, then $E$ belongs to $\langle\mathbb{R}\pi_{\ast}D^{b}_{\operatorname{coh}}(Y)\rangle_{n}$.

### v5

@@ 2. Generation
    $E$ is supported on $Z$ whenever $\operatorname{Supp}(E)$ is contained in $Z$; we say $E$ has full support if $\operatorname{Supp}(E)=X$.
3. Results
This section will prove our main results. To start, we need something that is akin to generic freeness of coherent $\mathcal{O}_{X}$-modules on a Noetherian integral scheme. The following might be known to a few, but we offer an alternative proof for the sake of interest in our setting.
Lemma 3.1.
Let $X$ be a Noetherian integral scheme. If $E$ and $G$ are objects of $D^{b}_{\operatorname{coh}}(X)$ with full support, then there exists an open immersion $i\colon U\to X$ such that $\mathbf{L}i^{\ast}E$ is in $\langle\mathbf{L}i^{\ast}G\rangle_{1}$ and $\mathbf{L}i^{\ast}G$ is in $\langle\mathbf{L}i^{\ast}E\rangle_{1}$.
Proof.
Consider a nonempty open affine $U$ in $X$. By [Let21, Proposition 3.5], the set $S$ of points $p$ in $U$ such that $E_{p}$ is in $\langle G_{p}\rangle_{1}$ is Zariski open. Clearly, $S$ is nonempty as it contains the generic point of $X$. By the same reasoning, the set $T$ of points $p$ in $X$ such that $G_{p}$ is in $\langle E_{p}\rangle_{1}$ is Zariski open and nonempty. Choose a nonempty open affine $W$ in $S\cap T$. Let $j\colon W\to X$ be the associated open immersion. From [Let21, Corollary 3.4], it follows that $\mathbf{L}j^{\ast}E$ is in $\langle\mathbf{L}j^{\ast}G\rangle_{1}$ and $\mathbf{L}j^{\ast}G$ is in $\langle\mathbf{L}j^{\ast}E\rangle_{1}$ because for each $p$ in $W$, one has $E_{p}$ is in $\langle G_{p}\rangle_{1}$ and $G_{p}$ is in $\langle E_{p}\rangle_{1}$. ∎
Proposition 3.2.
Let $X$ be a Noetherian scheme. Suppose $P\in\operatorname{Perf}(X)$ has full support. Then $D_{\operatorname{coh}}^{b}(X)$ is classically generated by $P$ and objects of the form $\mathbf{R}i_{\ast}\mathbf{L}i^{\ast}P$ where $i\colon Z\to X$ is a closed immersion from an integral scheme.
Proof.
Let $\mathcal{S}$ be the collection consisting of $P$ and objects of the form $\mathbf{R}i_{\ast}\mathbf{L}i^{\ast}P$ where $i\colon Z\to X$ is a closed immersion from an integral scheme. If $X=\emptyset$, then there is nothing to check, so without loss of generality, $X$ is nonempty. We argue by Noetherian induction. From [Lan24, Remark 2.23], we may impose that $X$ is reduced and the desired claim holds for any properly contained closed subscheme $Z$ of $X$. Let $E$ be an object of $D^{b}_{\operatorname{coh}}(X)$.
If $E$ has support properly contained in $X$, then the induction hypothesis tells us that $E$ belongs to $\mathcal{S}$. Indeed, let $Z$ be the reduced induced closed subscheme structure on $\operatorname{supp}(E)$ with corresponding ideal sheaf $\mathcal{I}$. By [Lan24, Remark 2.23], there is an integer $n$ such that $\mathcal{I}^{n}E=(0)$, which ensures $E$ is isomorphic to an object of the form $\mathbf{R}i_{\ast}E^{\prime}$ for $E^{\prime}$ in $D^{b}_{\operatorname{coh}}(Z_{n})$ and $i\colon Z_{n}\to X$ the closed immersion associated to the ideal sheaf $\mathcal{I}^{n}$. From [Tho97, Lemma 3.3], we know $\mathbf{L}i^{\ast}P\in\operatorname{Perf}(Z_{n})$ has full support. The induction hypothesis ensures that $\mathbf{R}i_{\ast}D^{b}_{\operatorname{coh}}(Z_{n})$ belongs to $\mathbf{R}i_{\ast}\mathcal{S}^{\prime}$ because $D^{b}_{\operatorname{coh}}(Z_{n})=\langle\mathcal{S}^{\prime}\rangle$ where $\mathcal{S}^{\prime}$ is the collection consisting of $\mathbf{L}i^{\ast}P$ and objects of the form $\mathbf{R}t_{\ast}\mathbf{L}(i\circ t)^{\ast}P$ where $t\colon Z^{\prime}\to Z_{n}$ is a closed immersion from an integral scheme. Note that $\mathbf{R}i_{\ast}\mathcal{S}^{\prime}$ is contained in $\mathcal{S}$, and so, it follows that $E$ belongs to $\langle\mathcal{S}\rangle$ as claimed.
We will now consider the case where $E$ has full support. Moreover, as objects in $D^{b}_{\operatorname{coh}}(X)$ are finitely built by their cohomology sheaves, we can impose that $E$ is a coherent $\mathcal{O}_{X}$-module. First, let us consider the case where $X$ is integral. We have that $E$ is a coherent $\mathcal{O}_{X}$-module with full support on the integral scheme $X$. By Lemma 3.1, there exists a dense open immersion $j\colon U\to X$ such that $\mathbf{L}j^{\ast}E\in\langle\mathbf{L}j^{\ast}P\rangle_{1}$ and $\mathbf{L}j^{\ast}G\in\langle\mathbf{L}j^{\ast}E\rangle_{1}$ in $D^{b}_{\operatorname{coh}}(U)$. If $U=X$, then we are done, so assume $U$ is a properly contained subset of $X$. Note that $\mathbf{L}j^{\ast}E$ is a direct summand of a bounded complex of the form $\oplus_{n\in\mathbb{Z}}\mathbf{L}j^{\ast}P^{\oplus r_{n}}[n]$, so we have an object $\mathbf{L}j^{\ast}A$ and an isomorphism of bounded complexes $\phi\colon\mathbf{L}j^{\ast}A\oplus\mathbf{L}j^{\ast}E\to\oplus_{n\in\mathbb{Z}}\mathbf{L}j^{\ast}P^{\oplus r_{n}}[n]$ in $D^{b}_{\operatorname{coh}}(U)$. There exists a roof in $D^{b}_{\operatorname{coh}}(X)$:
[…]
Observe that the cones of $f,g$ are supported on the properly contained closed subset $Z$ of $X$ as the map $\mathbf{L}j^{\ast}\colon D^{b}_{\operatorname{coh}}(X)\to D^{b}_{\operatorname{coh}}(U)$ is a Verdier localization. This ensures the objects $\operatorname{cone}(f),\operatorname{cone}(g)$ are in $\mathcal{S}$ by our inductive hypothesis. Therefore, $B$ belongs to $\mathcal{S}$, and hence, so does $A\oplus E$.
Lastly, we will consider the case where $X$ is not integral. Denote the irreducible components of $X=\operatorname{Supp}(E)$ by $Z_{1},\ldots,Z_{n}$. From [Sta23, Tag 01YD], there exists a short exact sequence of coherent $\mathcal{O}_{X}$-modules:
$0\to E_{1}\to E\to E_{1}^{\prime}\to 0$
where $\operatorname{Supp}(E_{1})\subseteq Z_{1}$ and $\operatorname{Supp}(E_{1}^{\prime})\subseteq\cup^{n}_{i=2}Z_{i}$. Both objects $E_{1}$ and $E_{1}^{\prime}$ are supported on properly contained closed subsets of $X$, and so, each belong to $\langle\mathcal{S}\rangle$ according to the induction hypothesis. However, this implies that $E$ belongs to $\langle\mathcal{S}\rangle$, which completes the proof. ∎
Theorem 3.3.
Let $X$ be a Noetherian scheme. Suppose $P$ is a classical generator for $\operatorname{Perf}(X)$. Then $D^{b}_{\operatorname{coh}}(X)$ is classically generated by $P$ and the objects of the form $\mathbf{R}i_{\ast}\mathbf{L}i^{\ast}P$ where $i\colon Z\to X$ is a closed immersion from an integral scheme such that $i(Z)$ is contained in $\operatorname{sing}(X)$.
Proof.
Denote by $\mathcal{U}$ for the collection of objects consisting of $P$ and those of the form $\mathbf{R}i_{\ast}\mathbf{L}i^{\ast}P$ where $i\colon Z\to X$ a closed immersion from an integral scheme such that $i(Z)$ is contained in $\operatorname{sing}(X)$. We will show that $\langle\mathcal{U}\rangle=D^{b}_{\operatorname{coh}}(X)$. By [Sta23, Tag 0BQT], we can see that $\mathcal{O}_{Z}$ is finitely built by $\mathbf{L}t^{\ast}P$ for each closed immersion $t\colon Z\to X$. So $\mathbf{R}t_{\ast}\mathcal{O}_{Z}$ is finitely built by $\mathbf{R}t_{\ast}\mathbf{L}t^{\ast}P$. This ensures that $\mathbf{R}i_{\ast}\mathcal{O}_{Z}$ is contained in $\langle\mathcal{U}\rangle$ for each closed immersion from an integral scheme such that $i(Z)$ is contained in $\operatorname{sing}(X)$.
Consider an affine open cover $X=\bigcup^{n}_{i=1}U_{i}$ with $s_{i}\colon U_{i}\to X$ being the associated open immersions. Note that for each $1\leq i\leq n$, one has $\langle\mathbf{L}s^{\ast}_{i}\mathcal{U}\rangle$ containing all objects of the form $p_{\ast}\mathcal{O}_{W}$ where $p\colon W\to U_{i}$ is a closed immersion where $W$ such that $p(W)$ is contained in $\operatorname{sing}(U_{i})$, including the structure sheaf. From [Tak14, Corollary 4.3.(1)], it follows that $\langle\mathbf{L}s^{\ast}_{i}\mathcal{U}\rangle=D^{b}_{\operatorname{coh}}(U_{i})$ because $U_{i}$ is affine for each $1\leq i\leq n$.
Let $E$ be an object of $D^{b}_{\operatorname{coh}}(X)$. There is an object $G_{i}$ in $\langle\mathcal{U}\rangle_{1}$ such that $\mathbf{L}s^{\ast}_{i}E$ is finitely built by $\mathbf{L}s^{\ast}_{i}G_{i}$ in $D^{b}_{\operatorname{coh}}(U_{i})$ because $\langle\mathbf{L}s^{\ast}_{i}\mathcal{U}\rangle=D^{b}_{\operatorname{coh}}(U_{i})$. Set $G=\oplus^{n}_{j=1}G_{i}$. It follows that $\mathbf{L}s^{\ast}_{j}E$ is finitely built by $\mathbf{L}s^{\ast}_{j}G$ in $D^{b}_{\operatorname{coh}}(U_{j})$ for each $1\leq j\leq n$. This ensures that $E_{p}$ is finitely built by $G_{p}$ in $D^{b}_{\operatorname{coh}}(\mathcal{O}_{X,p})$ for each $p$ in $X$. Hence, $E_{p}\overset{\mathbf{L}}{\otimes}K(p)$ is finitely built by $G_{p}$ in $D^{b}_{\operatorname{coh}}(\mathcal{O}_{X,p})$ for each $p$ in $X$ where $K(p)$ is the Koszul complex on a minimal set of generators for the maximal ideal of the local ring $\mathcal{O}_{X,p}$. Indeed, $K(p)$ is finitely built by $\mathcal{O}_{X,p}$, and tensoring with $E_{p}$ tells us that $E_{p}\overset{\mathbf{L}}{\otimes}K(p)$ is finitely built by $E_{p}$.
By [BIL+23, Theorem 1.7], if $E_{p}\overset{\mathbf{L}}{\otimes}K(p)$ is finitely built by $G_{p}$ in $D^{b}_{\operatorname{coh}}(\mathcal{O}_{X,p})$ for each $p$ in $X$, then $E$ is finitely built by $P\overset{\mathbf{L}}{\otimes}G$ in $D^{b}_{\operatorname{coh}}(X)$. As $G$ belongs to $\langle\mathcal{U}\rangle_{1}$, it is a direct summand of an object of the form $\oplus_{l\in\mathbb{Z}}(\mathbf{R}i_{l,\ast}\mathbf{L}i^{\ast}_{l}P)^{\oplus r_{l}}[l]$. By tensoring with $P$ and using projection formula, it follows that $P\overset{\mathbf{L}}{\otimes}\big{(}\oplus_{l\in\mathbb{Z}}(\mathbf{R}i_{l,\ast}\mathbf{L}i^{\ast}_{l}P)^{\oplus r_{l}}[l])$ is isomorphic to $\oplus_{l\in\mathbb{Z}}(\mathbf{R}i_{l,\ast}\mathbf{L}i^{\ast}_{l}(P\overset{\mathbf{L}}{\otimes}P))^{\oplus r_{l}}[l]$. Note that $P\overset{\mathbf{L}}{\otimes}G$ is a direct summand of $\oplus_{l\in\mathbb{Z}}(\mathbf{R}i_{l,\ast}\mathbf{L}i^{\ast}_{l}(P\overset{\mathbf{L}}{\otimes}P))^{\oplus r_{l}}[l]$. As $P\overset{\mathbf{L}}{\otimes}P$ is a classical generator for $\operatorname{Perf}(X)$, we see that $\oplus_{l\in\mathbb{Z}}(\mathbf{R}i_{l,\ast}\mathbf{L}i^{\ast}_{l}(P\overset{\mathbf{L}}{\otimes}P))^{\oplus r_{l}}[l]$ is finitely built by $\oplus_{l\in\mathbb{Z}}(\mathbf{R}i_{l,\ast}\mathbf{L}i^{\ast}_{l}P)^{\oplus r_{l}}[l]$. Therefore, $E$ must be contained in $\langle\mathcal{U}\rangle$. ∎
Remark 3.4.
It is important to assume $P$ is a classical generator for $\operatorname{Perf}(X)$ in Theorem 3.3. A possible relaxation might be imposing $P$ be a perfect complex on $X$ with full support, but there are counterexamples. For example, $P=\mathcal{O}_{X}$ where $X=\mathbb{P}^{1}_{k}$ for $k$ a field because $\mathcal{O}_{X}$ is not a strong generator for $D^{b}_{\operatorname{coh}}(\mathbb{P}^{1}_{k})$ and the singular locus is empty.
A useful application of Theorem 3.3 is when the singular locus is finite.
Example 3.5.
Let $X$ be an integral Noetherian scheme with at worst finitely many isolated singularities. Then the singular locus of $X$ is finite, say with closed points $p_{1},\ldots,p_{n}$ and associated closed immersions $\pi_{i}\colon\operatorname{Spec}(\kappa(p_{i}))\to X$. Suppose $P\in\operatorname{Perf}(X)$ is a classical generator for $\operatorname{Perf}(X)$. Then
$P\oplus(\oplus^{n}_{i=1}\pi_{i,\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p_{i}))})$
is a classical generator for $D^{b}_{\operatorname{coh}}(X)$. There are two cases where $P$ can be made more precise:
@@ 3. Results
    If $X$ is a quasi-projective variety over a field with very ample line bundle $\mathcal{L}$, then one can take $P=\oplus^{\dim X}_{i=0}\mathcal{L}^{\otimes i}$ (see e.g. [Orl09, Theorem 4]).
  2. (2)
    Suppose $X$ is a scheme that is projective over a Noetherian ring $R$. Denote the closed immersion by $t\colon X\to\mathbb{P}^{N}_{R}$. Set $\mathcal{L}$ to be the line bundle $t^{\ast}\mathcal{O}_{\mathbb{P}^{n}_{R}}(-1)$. By [Sta23, Tag 0A9V], we know that $\oplus^{N}_{i=0}\mathcal{O}_{\mathbb{P}^{n}_{R}}(-1)$ is a classical generator for $\operatorname{Perf}(\mathbb{P}^{n}_{R})$. However, the closed immersion $i$ is an affine morphism, so [Sta23, Tag0BQT] tells us that $i^{\ast}(\oplus^{N}_{i=0}\mathcal{O}_{\mathbb{P}^{n}_{R}}(-1))$ is a classical generator for $\operatorname{Perf}(X)$.
Now we work towards the rest. Initially, we showed that $\langle\mathbf{R}\pi_{\ast}D^{b}_{\operatorname{coh}}(Y)\rangle=D^{b}_{\operatorname{coh}}(X)$ for $\pi\colon Y\to X$ a proper surjective morphism by Noetherian induction coupled with Theorem 3.3. But we later realized how to refine this.
@@
Let $\pi\colon Y\to X$ be a proper surjective morphism of Noetherian schemes. For every object $K$ in $\mathbf{R}\pi_{\ast}D^{-}_{\operatorname{coh}}(Y)$ and integer $a$, there is a map $f\colon K\to L$ with $L$ in $\mathbf{R}\pi_{\ast}D^{b}_{\operatorname{coh}}(Y)$ such that $\operatorname{cone}(f)$ is in $D^{\leq a}_{\operatorname{Qcoh}}(X)$.
@@
Let $K:=\mathbf{R}\pi_{\ast}E$. Then $L=\mathbf{R}\pi_{\ast}(\tau_{\geq a-N}E)$ fits the bill for some integer $N$. ∎
@@
Let $\pi\colon Y\to X$ be a proper surjective morphism of Noetherian schemes. Let $n\geq 0$. For each $K$ in $D^{-}_{\operatorname{coh}}(X)$ which belongs to $\big{(}\mathbf{R}\pi_{\ast}D^{-}_{\operatorname{coh}}(Y)\big{)}^{\star n}$ and integer $a$, there is a map $f\colon K\to L$ with $L$ an object of $\big{(}\mathbf{R}\pi_{\ast}D^{b}_{\operatorname{coh}}(Y)\big{)}^{\star n}$ such that $\operatorname{cone}(f)$ in $D^{<a}_{\operatorname{Qcoh}}(X)$.
@@
where $A,B$ are objects of $\mathbf{R}\pi_{\ast}D^{-}_{\operatorname{coh}}(Y)$. Then Lemma 3.6 ensures there is a commutative square:
$A{{A^{\prime}}}B{{B^{\prime}}}$
@@
Let $\pi\colon Y\to X$ be a proper surjective morphism of Noetherian schemes. If $E$ is an object of $D^{b}_{\operatorname{coh}}(X)\cap\langle\mathbf{R}\pi_{\ast}D^{-}_{\operatorname{coh}}(Y)\rangle_{n}$, then $E$ belongs to $\langle\mathbf{R}\pi_{\ast}D^{b}_{\operatorname{coh}}(Y)\rangle_{n}$.